### v1

# Learning Online Belief Prediction for Efficient POMDP Planning in Autonomous Driving

Abstract
Effective decision-making in autonomous driving relies on accurate inference of other traffic agents’ future behaviors. To achieve this, we propose an online learning-based behavior prediction model and an efficient planner for Partially Observable Markov Decision Processes (POMDPs). We develop a Transformer-based prediction model, enhanced with a recurrent neural memory model, to dynamically update latent belief states and infer the intentions of other agents. The model can also integrate the ego vehicle’s intentions to reflect closed-loop interactions among agents, and it learns from both offline data and online interactions. For planning, we employ an option-based Monte-Carlo Tree Search (MCTS) planner, which reduces computational complexity by searching over action sequences. Inside the MCTS planner, we use predicted long-term multi-modal trajectories to approximate future updates, which eliminates iterative belief updating and improves the running efficiency. Our approach also incorporates deep Q-learning (DQN) as a search prior, which significantly improves the performance of the MCTS planner. Experimental results from simulated environments validate the effectiveness of our proposed method. The online belief update model can significantly enhance the accuracy and temporal consistency of predictions, leading to improved decision-making performance. Employing DQN as a search prior in the MCTS planner considerably boosts its performance and outperforms an imitation learning-based prior. Additionally, we show that the option-based MCTS substantially outperforms the vanilla method in terms of performance and efficiency.
Index Terms:
POMDP, Behavior Prediction, Belief Update, Monte-Carlo Tree Search, Interactive planning
I Introduction
Decision-making under uncertainties is crucial for the safety of autonomous driving systems. In particular, the uncertainty in human road users’ behaviors is a primary source of uncertainties [1], which imposes significant challenges to the safe navigation of autonomous vehicles (AVs) in real-world scenarios. The Partially Observable Markov Decision Process (POMDP) [2, 3] offers a mathematically sound framework to address this problem. However, current POMDP planners have limited capabilities, as they only represent the hidden states of other agents with specific semantic meanings and cannot integrate with deep learning-based prediction models. This restricts their capability and scalability in more diverse and complex real-world situations. Therefore, we aim to develop a deep learning-based online behavior prediction model and an efficient planner for POMDPs in autonomous driving.
Figure 1: Illustration of our proposed planning approach. We utilize a neural memory-based belief update model to continually update the human agent’s intentions over time based on new observations and the AV’s actions. An option-based MCTS planner, guided by a learned Q-value function, searches for the optimal action based on current belief states.
One challenge involves adapting an existing open-loop prediction model to an online and closed-loop setting. We propose a neural memory-based belief update model that optimizes closed-loop prediction performance and learns through online interactions. Specifically, our proposed model includes a Transformer-based encoder that maps the observation of an agent (in the context of maps and other agents) to latent space. At each time step in online training or testing, we utilize a gated recurrent unit (GRU) model to update the latent belief state of each tracked agent by considering its last latent belief state, the current latent observation, and the intention of the ego agent. We then employ a decoder that maps the latent belief state back into the explicit belief state, represented by a probabilistic distribution of intentions (long-term trajectories) for downstream planning. Our model can significantly improve the consistency and probability estimation compared to offline prediction models that are designed, trained, and evaluated in an open-loop manner. Our model is also able to capture the dynamic interactions between the AV and other agents, as opposed to open-loop conditional prediction models [4, 5, 6].
Another challenge lies in developing a computationally efficient planner from POMDPs. To address this, we adopt the Monte-Carlo tree search (MCTS) algorithm [7, 8, 9] and incorporate several enhancements to reduce the computational complexity for real-world applicability. First, we leverage an option-based MCTS method that searches over options (i.e., action sequences or motion primitives) [10, 11], allowing a more in-depth search within a limited computation budget. Additionally, we use predicted multi-modal long-term trajectories from the online prediction model to approximate future updates for other agents within the planning horizon, which improves computational efficiency while ensuring planning performance. Lastly, we train a deep Q-value network based on the environment reward feedback during the online learning process, which is used as the search prior for the MCTS planner and significantly improves its planning performance.
The core idea of our approach is illustrated in Fig. 1, and the main contributions of this paper are summarized as follows.
  1. 1.
    We propose a neural memory-based belief update model that aids in online and closed-loop behavior prediction for POMDP planning.
  2. 2.
    We introduce an option-based MCTS planning method, coupled with a Q-value function network that acts as a heuristic guide during the search process.
  3. 3.
    We establish an online learning framework of the belief prediction and Q-value network and validate our proposed method with a large-scale driving dataset and simulated driving environment.
II Related Work
II-A Behavior Prediction
Recent advancements in deep learning-based prediction models have greatly improved accuracy in various benchmarks [12, 13, 14]. However, when applying these models in online and closed-loop settings, temporal consistency becomes a major problem because these models are all designed to work with independent frames and are trained offline. To address this problem, the motion forecasting with self-consistent constraints (MISC) method [15] has proposed self-consistent losses in training. However, MISC still uses independent frames as model inputs and is trained offline. The concept of streaming motion prediction [16] has been proposed recently, where a neural differentiable filter is used to update the predicted states of agents at each timestep. Our method differs from it in that we update the latent belief states of agents, which ensures temporal consistency, improves prediction accuracy, and allows for incorporating the AV’s influence in the update. Moreover, many studies focus on ego-conditioned prediction (i.e., incorporating the future actions of the AV into the predictions of other agents) to make interaction-aware decisions [17, 18, 19]. However, most of these models are trained in an open-loop manner. In contrast, our model is trained online and capable of modeling closed-loop interactions.
II-B POMDP in Autonomous Driving
POMDP is a mathematically principled framework for decision-making under uncertainty, which has been applied in autonomous driving tasks. [20] proposed a POMDP planner for automated driving that considers human internal states, represented by values of the parameters of driving behavior models. [1] used possible routes followed by other vehicles as the hidden states in POMDP, and the belief states can be estimated over time by particle filter. [21] proposed a POMDP planning method by estimating drivers’ intelligence levels and degrees of rationality through probabilistic inference. However, these planners only infer the semantic parameters of some dynamics or probabilistic models, which limits their capability in more complex scenarios with heterogeneous agents. In contrast, our approach employs latent belief states for other agents, which can be recovered to their future trajectories. Another thread of research is to incorporate neural network models in POMDP [22, 23, 24], and one effective approach is recurrent latent (belief) state update [25]. Therefore, we use the recurrent GRU model to track the latent belief states of other agents and iteratively update their belief states while considering the intention of the ego vehicle.
II-C MCTS Planning in Autonomous Driving
MCTS algorithm has achieved great success in solving complex MDP problems such as board games and Atari games [26], and it has been increasingly applied to autonomous driving tasks [7, 8, 27]. The MuZero method [28] has significantly improved the scalability of MCTS by integrating learned models, such as transition dynamics and value functions. Our work focuses on learning the behavior prediction model for other road users to approximate future transitions. Additionally, to overcome the issue of high-dimensional and continuous search space, one practical solution is using an option-based approach that searches over action sequences instead of raw control actions [10], which effectively reduces search depth. Another improvement is to use learned heuristic models to guide the tree search process and find the optimal solution. For instance, [7] and [8] incorporated a trajectory prediction model as heuristics in tree search, improving both efficiency and performance. In our work, we utilize deep Q-learning as a search prior in the MCTS planner, which shows significant improvements in planning performance.
Figure 2: Illustration of the proposed POMDP decision-making framework. (a) Diagram of the POMDP planning process. At each time step, we use a Transformer encoder to map the observation into latent space, which is used to update the latent belief state. We use an MLP decoder to project the latent belief state into probabilistic future trajectories. (b) Option-based MCTS planner. The planner takes as input the estimated future states of other agents and searches over options. Additionally, we employ a learned Q-value network to guide the selection step at the root node. (c) Structure of the belief update network. The information of the previous latent belief state, current latent observation, and the ego agent’s action are fused in the update process.
III Method
III-A Problem Statement
We formulate the decision-making problem as a POMDP, which is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{Z},\mathcal{T},\mathcal{R})$. $\mathcal{S}$ is the state space and $\mathbf{s}_{t}^{i}$ denotes the state of agent $i$ and time $t$. $\mathbf{a}_{t}^{e}\in\mathcal{A}$ is the action space for the ego agent. $[\mathbf{o}_{t}^{i},\mathbf{o}_{t}^{m}]\in\mathcal{O}$ is the observation space, where $\mathbf{o}^{i}_{t}$ denotes the observed physical state of the agent (e.g., position, orientation, velocity, and size) and $\mathbf{o}^{m}_{t}$ is the shared road map. $\mathcal{Z}(\mathbf{s}_{t}|\mathbf{o}_{t})$ is the observation function, $\mathcal{T}(\mathbf{s}_{t+1}|\mathbf{a}_{t},\mathbf{s}_{t})$ is the state transition function, and $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})$ is the immediate reward function.
The ego (decision-making) agent cannot directly observe the internal state $\mathbf{s}^{i}$ of other agents, which can encompass attributes such as the agent’s navigational goals, behavioral traits, and intentions [2]. Therefore, the ego agent maintains a belief $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ as an estimate of the states of its surrounding $N$ agents. The belief state starts with an initial belief $\mathbf{b}_{0}^{i}$ and is updated $\mathbf{b}_{t}^{i}$ at each time step $t$ according to a belief update function $\rho$:
$\mathbf{b}^{i}_{t}=\rho(\mathbf{b}^{i}_{t-1},\mathbf{a}^{e}_{t-1},\mathbf{o}_{t}^{i}),$ (1)
where $\mathbf{a}^{e}$ is the action of the ego agent.
We assume that the ego agent makes decisions based on its current belief of the surrounding agents; thus, the entire history of observations is taken into consideration implicitly in the ego agent’s actions. We aim to find the optimal ego policy, denoted by $\pi(\mathbf{a}_{t}|\mathbf{b}_{t})$, that maximizes the expected total discounted reward:
$\pi^{*}=\arg\max_{\pi}\ \mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}\left(\mathbf{s}_{t},\pi(\mathbf{b}_{t})\right)\right],$ (2)
where $T$ is the horizon and $\gamma\in(0,1)$ is the discount factor.
Specifically, we represent the state of an agent $\mathbf{s}^{i}$ as its intention or planned future trajectory. We maintain a latent belief state of each agent in our model, which can be decoded back to an explicit belief state, such as probabilistic intentions or multi-modal future trajectories. To achieve this, we employ an encoder network that maps the observation to a latent space, and a GRU network to update the latent belief state for each agent, which is then decoded to possible trajectories. The overall POMDP planning framework is illustrated in Fig. 2.
III-B Neural Recurrent Belief Update
Observation Encoding. The observation space includes the historical trajectories of traffic agents (including the ego) $\mathbf{o}^{a}\in\mathbb{R}^{N_{a}\times T_{h}\times D_{a}}$ and map polylines $\mathbf{o}^{m}\in\mathbb{R}^{N_{m}\times N_{w}\times D_{m}}$, where $N_{a}$ is the number of agents, $T_{f}$ is the fixed history steps, $N_{m}$ is the number of polylines, and $N_{w}$ is the number of waypoints in a polyline. Note that their positional attributes are normalized according to their respective local coordinates. We encode the historical states of agents using GRU networks and multi-layer perceptron (MLP) networks to encode the map elements, and we apply max-pooling to reduce the waypoint axis. Finally, we concatenate these elements to obtain an encoding of the scene $\mathbf{h}\in\mathbb{R}^{(N_{a}+N_{m})\times D}$, where $D$ is the hidden feature dimension.
Subsequently, we employ a Transformer network with query-centric multi-head attention [12, 14] to extract the relationships between these scene elements in a symmetric manner. Since each agent is encoded in their local coordinates rather than a global coordinate, it is invariant to the position change of the ego agent and thus facilitates belief update. Specifically, we apply pair-wise attention calculation for each query token from encoding $\mathbf{h}$, and the query, key, and value inputs of the attention module are derived as follows.
$\begin{split}Q^{i}&=\mathbf{h}^{i},\ K^{ij}=\mathbf{h}^{j}+\mathbf{e}^{ij},\ V^{ij}=\mathbf{h}^{j}+\mathbf{e}^{ij},\\ \mathbf{e}^{ij}&=PE(\Delta_{x}^{ij},\Delta_{y}^{ij},\Delta_{heading}^{ij}),\end{split}$ (3)
where $i,j$ are the indexes of query tokens, and $PE$ stands for positional encoding, which is a feed-forward network.
We denote the encoding process as a function with parameters $f_{\theta}$ and the final encoding as $\mathbf{z}\in\mathbb{R}^{(N_{a}+N_{m})\times D}$, and we can retrieve $N$ agents of interest from the scene encoding.
Belief Update. We denote the belief update network as $f_{\rho}$, which takes as input the last latent belief of the agent $\mathbf{\tilde{b}}_{t-1}^{i}\in\mathbb{R}^{D}$, the current latent observation $\mathbf{z}_{t}^{i}$, as well as the ego agent’s selected option (short-term planned trajectory) relative to the agent $\mathbf{a}_{t-1}^{i,e}\in\mathbb{R}^{T_{o}\times 3}$. We set the initial latent belief as $\mathbf{\tilde{b}}_{0}^{i}=\mathbf{z}_{0}^{i}$, and the belief update process can be formulated as follows.
$\begin{split}\mathbf{z}_{t}^{i}&=f_{\theta}(\mathbf{o}_{t}^{i},\mathbf{o}_{t}^{\neg i},\mathbf{o}_{t}^{m}),\\ \mathbf{\tilde{b}}_{t}^{i}&=f_{\rho}(\mathbf{z}_{t}^{i},\mathbf{\tilde{b}}_{t-1}^{i},\mathbf{a}_{t-1}^{i,e}),\\ \mathbf{b}^{i}_{t}&=f_{\phi}(\mathbf{\tilde{b}}_{t}^{i}),\end{split}$ (4)
where $\mathbf{b}^{i}_{t}$ represents the explicit belief state, and $f_{\phi}$ is the trajectory decoder network that decodes multimodal and uncertainty predictions from the latent belief state.
The structure of the belief update network is illustrated in Fig. 2(c). We first account for the influence of the ego agent’s action on the belief of the tracked agent. For this, we use two MLPs - one is utilized to encode the ego agent’s relative trajectory, and the other is used to predict a gating value. We use this gating value to control the flow of information to the belief update because some agents may not be influenced by the ego agent. Afterward, the ego agent’s influence is combined with the latent observation to update the belief state through the GRU cell. Note that some agents may be occluded for a few timesteps, and we take the last available belief state of those agents in the update.
Trajectory Decoding. Given an agent’s latent belief state $\mathbf{\tilde{b}}_{t}^{i}$, we use an MLP $f_{\phi}$ to decode from the latent belief to explicit belief, which is represented by a Gaussian mixture model (GMM) of future trajectories $\mathbf{b}^{i}\in\mathbb{R}^{M\times T_{f}\times D_{f}}$. This GMM comprises $M$ modalities and $T_{f}$ future timesteps, and features at each step include $x$ and $y$ coordinates, variances $\sigma_{x}$ and $\sigma_{y}$, and probability $p$. It is important to note that the decoder $f_{\phi}$ can also be used for the latent observation $\mathbf{z}^{i}$, as in most existing open-loop motion prediction models.
III-C Option-based MCTS Planner
We use the belief states of agents $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ (multi-modal long-term trajectories) to approximate the future updates within the planning horizon. This effectively simplifies the POMDP problem into an MDP problem [20], which can be efficiently solved using the MCTS algorithm. We employ receding horizon planning and construct a new policy for the current belief each time a new observation is received. Furthermore, to optimize the quality and efficiency of the MCTS planner, we employ an option-based approach.
Options. An option is a sequence of actions over a time duration [11], and options can be learned from data and can differ in length [10, 29]. For simplicity, we employ fixed-length heuristic options (high-level actions, like overtaking and turning). We assume that the routing module provides a reference path, such that we can generate options based on this reference path. At each node, options are generated using the available actions for acceleration ($a_{s}$) and lateral speed ($v_{l}$), and the resulting states are calculated in the Frenet frame:
$\footnotesize\begin{split}&\ddot{s}(\tau)=a_{s},\ \dot{s}(\tau)=\dot{s}(0)+a_{s}\tau,\ s(\tau)=s(0)+\dot{s}(0)\tau+\frac{1}{2}a_{s}\tau^{2},\\ &\dot{l}(\tau)=v_{l},\ l(\tau)=l(0)+v_{l}\tau,\ \tau\in[1,2,\cdots,T_{o}]\Delta\tau,\end{split}$ (5)
where $s$ and $l$ represent the longitudinal and lateral directions along the reference path, $s(0)$, $l(0)$, $\dot{s}(0)$ are the initial state of the node, $\Delta\tau$ is the time interval, and $T_{o}$ is the option length. Then, we can transform the states in the Frenet frame back into the Cartesian frame $\left(x(\tau),y(\tau),v(\tau),a(\tau)\right)$ [30].
Objective Function. We calculate the cost of options considering collision risk, ride comfort, deviation from the route, and desired speed. The objective function of the planner is:
$\max\sum_{n=0}^{T_{d}-1}\gamma^{nT_{o}\Delta\tau}\sum_{i}-w_{i}c_{i}(n),$ (6)
where $T_{d}=\frac{T_{f}}{T_{o}}$ is the search depth, $w_{i}$ is the weight, and $c_{i}$ is the cost term at time segment $n$. The ride comfort factor includes acceleration, jerk, and lateral acceleration of the trajectory, and the desired speed cost is the difference between the current speed and speed limit. Each cost term is averaged across all timesteps in the node.
Importantly, the collision risk of an option is calculated as:
$c_{r}(n)=\sum_{i}^{N}\sum_{j}^{M}\mathds{1}(p_{j}\geq p_{th})p_{j}\mathds{1}_{overlap}\left(\mathbf{s}^{e}(n),\mathbf{b}^{i}(j,n)\right),$ (7)
where $\mathbf{s}^{e}$ is the state of the ego agent determined by the generated options, $\mathbf{b}^{i}(j)$ is the $j$-th predicted trajectory of agent $i$, $p_{j}$ is the predicted likelihood of that trajectory. $\mathds{1}_{overlap}$ is a binary indicator to check whether the two trajectories collide, and $p_{th}$ is a threshold to filter out low probability futures.
Algorithm 1 MCTS with online belief prediction model
1:MCTS planner $TP$, observation encoder $f_{\theta}$, belief update network $f_{\rho}$, trajectory decoder $f_{\phi}$, Q-value network $Q_{\epsilon}$, $N$ number of predicted agents.
2:Initialize timestep $t\leftarrow 0$ and belief state tracker $B\leftarrow\O$
3:while task is active do
4:     Observe environment $\mathbf{o}_{t}=\left[\{\mathbf{o}_{t}^{i}\}_{i=1:N_{a}},\mathbf{o}^{m}_{t}\right]$
5:     Encode observation and retrieve $N$ agent tokens
6:      of interest $\{\mathbf{z}_{t}^{i}\}_{i=1:N}=f_{\theta}(\mathbf{o}_{t})$
7:     for $i\leftarrow 1$ to $N$ do
8:         if $\mathbf{\tilde{b}}_{t-1}^{i}\in B$ then
9:              $\mathbf{\tilde{b}}_{t}^{i}=f_{\rho}(\mathbf{z}_{t}^{i},\mathbf{\tilde{b}}_{t-1}^{i},\mathbf{a}_{t-1}^{i,e})$// Update belief state
10:         else
11:              $\mathbf{\tilde{b}}_{t}^{i}=\mathbf{z}_{t}^{i}$ // Initialize belief state
12:         end if
13:         Update state tracker $B\leftarrow B\cup\{\mathbf{\tilde{b}}_{t}^{i}\}$
14:         Decode probabilistic trajectories $\mathbf{b}_{t}^{i}=f_{\phi}(\mathbf{\tilde{b}}_{t}^{i})$
15:     end for
16:     Calculate prior $Pr(\mathbf{z}_{t}^{e},\mathbf{a}^{e})$ for root node using Eq. 9
17:     Compute action sequence ${\mathbf{a}}^{*}\leftarrow TP(\{\mathbf{b}_{t}^{i}\}_{i=1:N},Pr)$
18:     Execute first action ${\mathbf{a}_{t}^{e}}\leftarrow\mathbf{a}_{0}^{*}$
19:     Increment timestep $t\leftarrow t+1$
20:end while
MCTS Algorithm. MCTS consists of several key steps: selection, expansion, evaluation, and backup, and this process is repeated multiple times until a termination condition is reached. We focus on the selection step in the following, and more details about MCTS can be found in [7]. In the selection step, the nodes are selected according to:
$a=\arg\max_{a\in A}\tilde{Q}(s,a)+Pr(a)C_{p}\sqrt{\frac{2\log N_{p}}{n_{a}+1}},$ (8)
where $\tilde{Q}(s,a)$ is the expected return, $n_{a}$ is the number of times the node has been visited, $N_{p}$ is the total number of times the parent node has been visited, and $C_{p}$ is a temperature parameter used to balance exploration and exploitation.
The prior $Pr(a)$ is given by the learned Q-value network $Q_{\epsilon}$, and we apply this prior to the root node, while setting the prior to a uniform distribution in other nodes. The sampling prior for the root node is calculated as:
$Pr(a_{i})=\frac{\exp Q_{\epsilon}(\mathbf{z}^{e}_{t},a_{i})}{\sum_{j}\exp Q_{\epsilon}(\mathbf{z}^{e}_{t},a_{j})},$ (9)
where $\mathbf{z}^{e}_{t}$ is the latent observation state for the ego agent at the current time, which is retrieved from the observation encoder and contains necessary information including its past states, other agents, and road maps.
The proposed POMDP deep predictive planning algorithm can be summarized in Algorithm 1.
III-D Learning Framework
Offline Learning. In this stage, we train the observation encoder $f_{\theta}$ with the multi-modal trajectory decoder $f_{\phi}$ using an offline driving dataset. We employ the following loss on the selected positive GMM component (the mode with the smallest L2 distance to ground truth) for a specific agent at a future time step.
$\footnotesize\mathcal{L}_{GMM}=\log\hat{\sigma}_{x}+\log\hat{\sigma}_{y}+\frac{1}{2}\left(\left(\frac{s_{x}-\hat{s}_{x}}{\hat{\sigma}_{x}}\right)^{2}+\left(\frac{s_{y}-\hat{s}_{y}}{\hat{\sigma}_{y}}\right)^{2}\right)\!-\log\hat{p},$ (10)
where ($\hat{s}_{x},\hat{s}_{y},\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{p}$) is the selected Gaussian component; $s_{x},s_{y}$ are the ground-truth future states of the predicted agent.
Algorithm 2 Online learning process
1:pre-trained observation encoder $f_{\theta}$, belief update network $f_{\rho}$, trajectory decoder $f_{\phi}$, twin Q-value network $Q_{\epsilon}^{1,2}$, training steps $N_{t}$, exploration rate $\lambda$
2:Initialize the replay buffer $R$
3:for $i\leftarrow 1$ to $N_{t}$ do
4:     Initialize episodic observation buffer $O$ and trajectory
5:      prediction buffer $P$
6:     while episode not terminated do
7:         Collect and store agent observations in $O$
8:         Encode observations to latent space $z$ using $f_{\theta}$
9:         Predict trajectories using $f_{\rho}$ and $f_{\phi}$ and store in $P$
10:         Select action $a$: with probability $\lambda$, sample random
11:            action; otherwise use MCTS planner
12:         Execute action $a$, observe reward $r$, and obtain
13:          next latent state $z^{\prime}$
14:         Store transition $(z,a,r,z^{\prime})$ in replay buffer $R$
15:         Update $Q_{\epsilon}^{1,2}$ by sampling from $R$
16:     end while
17:     Update $f_{\rho}$ and $f_{\phi}$ using prediction results from $P$ and
18:      ground truth data from $O$ in hindsight
19:end for
20:return Trained networks $f_{\rho}$, $f_{\phi}$, $Q_{\epsilon}^{1}$
Online Learning. In this stage, we fix the weights of the observation encoder, and we learn the belief update model $f_{\rho}$ and fine-tune the trajectory decoder $f_{\phi}$ in a driving simulator. At the end of an episode, we obtain a prediction buffer and an observation buffer that contains the positions of all agents at each timestep. We utilize the GMM loss in Eq. 10 to minimize the prediction loss at each valid time step in hindsight. In addition, we leverage the clipped double Q-learning [31] to learn a Q-value network $Q_{\epsilon}$ to guide the tree search planner. The reward function from the environment for Q-learning is formulated as:
$r=r_{col}+r_{prog}-0.01r_{expert},$ (11)
where $r_{col}=-10$ is the penalty imposed if the ego vehicle collides with any other agents, $r_{prog}$ is the distance traveled by the ego vehicle, and $r_{expert}$ is the state difference between the agent and logged human expert, which could encode other driving features that are hard to manually design. Since the true state of the environment can only be obtained at the current step, and future states are approximated by the model, the Q-value network is used to guide the selection at the root node.
We jointly learn the belief update and Q-network (with the MCTS planner) to improve the overall decision-making performance. The procedure for online belief update learning and Q-learning is summarized in Algorithm 2.
IV Experiments
IV-A Dataset and Simulator
We employ the Waymo Open Motion Dataset (WOMD) [32] to train the models and replay the traffic data for surrounding agents in the MetaDrive [33] simulator to conduct online training and testing. Additionally, we set the agent behavior to be reactive in the simulator, allowing other agents to perform basic reactive actions, such as avoiding collisions from the rear end of the ego vehicle. Fig. 3 shows some replayed scenarios from WOMD in the MetaDrive simulator and the obtained observation space for the ego agent.
Figure 3: Examples of replayed WOMD scenarios in the MetaDrive simulator and the processed observation space including surrounding map polylines and historical trajectories of agents.
IV-B Implementation Details
Data. In the offline learning stage, we select 20,000 scenarios from the WOMD, each with a duration of 20 seconds. For online learning, we randomly choose 2,000 scenarios from the training set for simulation-based replay. In the testing phase, we select an additional 200 scenarios from the WOMD. Note that we exclude static scenarios where the ego vehicle travels less than 20 meters, such as waiting at red lights.
Neural Networks. We select the closest $N_{a}=16$ agents to the ego vehicle and their observed historical horizon is $T_{h}=20$ timesteps. Additionally, we incorporate $N_{m}=50$ map polylines in encoding, each comprising $N_{w}=20$ waypoints. The Transformer encoder has $3$ attention layers and a hidden dimension of $D=256$. The trajectory decoder outputs $M=6$ possible trajectories for an agent, $8$ seconds into the future $T_{f}=80$. In the belief update network, the ego agent’s action is represented by a short-term trajectory ($T_{o}=20$), which is processed relative to each tracked agent and encoded. The Q-value network is an MLP that takes as input the hidden state of the ego agent and outputs the values of actions.
MCTS Planner. The planning horizon is set to $T_{f}=80$ timesteps, with a time interval of $\Delta t=0.1\ s$, and the option length is set to $T_{o}=20$. This significantly reduces the search depth to $T_{d}=4$. We track the belief states of $N=8$ agents and set the probability threshold for calculating collision risk to $p_{th}=0.15$. The options at each node are derived from combinations of lateral velocity $v_{l}=[-1,0,1]\ m/s$ and longitudinal acceleration $a=[-4,-2,0,1,3]\ m/s^{2}$. Lateral velocity is only considered if the acceleration is $-2,0,1$, leading to a total of $11$ options. We also ensure that the velocity remains non-negative. The planner executes a total of $100$ search iterations and applies a discount factor of $\gamma=0.8$.
Offline learning. We utilize the AdamW optimizer for model training, conducting $40$ training epochs. The learning rate begins at $42e-$ 4 and is halved every $5$ epochs. The model is trained on an NVIDIA A100 GPU with a batch size of 256.
Online Learning. We train the models in the simulator with $N_{t}=300,000$ steps. We employ Adam optimizer to train the networks and the learning rate starts with $43e-$ 4 and reduces by $30\%$ every $50,000$ steps. The exploration rate $\lambda$ starts with $0.8$ and decays to $0.05$ in the middle of training. The discount factor for Q-learning is $\gamma=0.99$, the replay buffer size is $100,000$, and the batch size is $128$.
Testing. The runtime of the prediction model and MCTS planner is tested on a laptop equipped with an AMD 7945HX CPU and NVIDIA RTX 4060 GPU.
IV-C Main Results
Baselines. In testing, we compare our method against three strong baseline approaches: RL (DQN) method [31], intelligent driver model (IDM) [34], and differentiable integrated prediction and planning (DIPP) [35]. In addition, we introduce several ablated versions of our approach. The offline model refers to the prediction model without online belief update, and it is only trained on the offline dataset. The online update model has two versions: one with the ego agent’s intentions in the online belief update, and one without. Furthermore, we evaluate the effectiveness of employing imitation learning (IL) (i.e., predicted future trajectory for the ego vehicle) [7, 8] and DQN as prior in the MCTS planner. In training, we introduce two baselines for comparison: DQN agent and MCTS planner with online belief update model but lacking Q-learning.
Metrics. The task-related metrics include success rate, average episodic reward, average episodic time (number of steps), and average log divergence (L2 distance) measuring the deviation of the ego vehicle from logged behavior. Regarding prediction performance, we employ three key metrics: minimum average displacement error (minADE), consistency (variation in predictions across consecutive frames), and score accuracy (whether the highest probability is correctly assigned to the closest modality to the ground truth). The prediction consistency is defined as:
$\frac{1}{N}\sum_{i=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\frac{1}{T_{f}-1}\sum_{\tau=1}^{T_{f}-1}\left\lVert\mathbf{b}^{i}_{t}(m,\tau)-\mathbf{b}^{i}_{t-1}(m,\tau+1)\right\rVert^{2},$
where $\mathbf{b}_{t-1}$ is the prediction result from the last time frame, and this metric is averaged across all timesteps in the episode.
Training Results. We train different methods using three random seeds, and the training results are depicted in Fig. 4. The DQN agent shows the worst performance in terms of success rate, and learning the online prediction model without Q-learning significantly influences the MCTS planner’s effectiveness, mainly due to the deviation from expert trajectories. While the MCTS planner with an online prediction model and deep Q-learning shows favorable effectiveness, its performance can be further enhanced with the online model that integrates the ego agent’s intentions into the belief update process. This approach achieves the highest reward, highlighting the efficacy of closed-loop interaction modeling.
Figure 4: Training results of our proposed method against baseline methods. Left: average episodic reward; right: average success rate.
TABLE I: Performance comparison of different decision-making approaches in the testing scenarios
Method Success rate Task time Log divergence Episode reward Prediction minADE Consistency Score accuracy
IDM Agent 0.91 132.18 11.80 38.59
DIPP [35] 0.94 125.65 10.35 43.88 0.902 1.631 0.260
DQN Agent 0.53$\pm$0.04 62.32$\pm$10.74 11.87$\pm$0.58 30.08$\pm$1.35
MCTS + Offline model 0.78$\pm$0.02 174.20$\pm$2.12 10.43$\pm$0.10 17.86$\pm$0.80 1.025 1.652 0.261
MCTS + IL + Offline model 0.72$\pm$0.04 175.83$\pm$1.37 10.89$\pm$0.16 16.84$\pm$1.17 1.019 1.674 0.264
MCTS + Online update 0.82$\pm$0.01 168.29$\pm$1.08 10.59$\pm$0.24 21.28$\pm$0.75 1.047 0.574 0.314
MCTS + IL + Online update 0.87$\pm$0.02 141.06$\pm$1.65 10.57$\pm$0.16 31.86$\pm$0.84 0.984 0.573 0.315
MCTS + DQN + Offline model 0.94$\pm$0.01 119.58$\pm$0.41 10.43$\pm$0.10 45.31$\pm$0.52 0.847 1.712 0.266
MCTS + DQN + Online update 0.98$\pm$0.02 122.24$\pm$1.23 10.35$\pm$0.07 46.21$\pm$0.18 0.829 0.570 0.320
MCTS + DQN + Online update (w/ ego) 0.98$\pm$0.02 119.50$\pm$0.94 10.27$\pm$0.12 46.62$\pm$0.26 0.878 0.562 0.323
Figure 5: Qualitative comparisons of different behavior prediction models and their influences on the decision-making task. For clarity, only the most likely modality, along with its predicted probability, is displayed, and the length of trajectories is 5 seconds. The offline prediction model shows significant fluctuations in trajectories and probabilities, which cause the AV to fail to move ahead of the interacting agent in Scenario 1. In contrast, our online update model shows consistent predictions over time, with the accuracy gradually increasing when receiving new observations. The enhancements in behavior prediction provided by our online update model enable the MCTS planner to make more human-like decisions.
Testing Results. We evaluate our proposed method against other baselines in the testing scenarios, and the results are summarized in Table I. The key conclusions and findings are as follows. (1) The DQN agent shows the weakest performance, while both the rule-based IDM agent and the optimization-based DIPP method demonstrate favorable efficacy in testing. (2) The basic MCTS method without DQN guidance performs significantly worse, and DQN guidance is more effective than IL guidance, particularly when IL guidance is learned offline. (3) The prediction consistency of an offline-learned model is substantially worse compared to that of online belief update models. While offering similar prediction accuracy, these online belief update models significantly enhance prediction consistency and the accuracy of intention probability estimations, leading to improved decision-making performance. (4) Incorporating the ego vehicle’s intention into the belief update process in the online prediction model can further improve the decision-making performance in terms of the episode reward. Some qualitative results are presented in Fig. 5, which reveal that our proposed online update model provides more consistent and accurate trajectory predictions of interacting agents compared to the offline prediction model. The online prediction model is able to continuously improve its prediction accuracy (in terms of trajectory and probability) over time by receiving new observations. See closed-loop testing results.
IV-D Ablation Study
Influence of option length. We investigate the impact of varying option lengths on the performance of the MCTS planner, as detailed in Table II. We find that shorter option lengths (1, 5, 10) lead to increased computation time and lower success rate and episode reward. This is because, within a limited number of search iterations, a deeper search depth makes it increasingly difficult to find the optimal action. Conversely, extending the option length to 40, which essentially simplifies the process into two-stage planning, yields a success rate comparable to an option length of 20, but the episode reward at this length is reduced. Consequently, an option length of 20 appears to provide a good balance between performance and computational efficiency.
TABLE II: Influence of option length on planning performance
Length Success rate Episode reward Log divergence Runtime ($ms$)
1 0.59 21.47 14.34 834.8
5 0.62 25.90 13.58 376.5
10 0.88 37.12 12.08 196.4
20 0.98 46.62 10.27 126.5
40 0.98 42.25 10.53 92.4
Influence of probability threshold. We examine the effects of the probability threshold in the collision risk cost (Eq. 7) on planning performance. Setting $p_{th}=0$ includes all predicted futures in planning, and we also set up an alternative approach that only considers the most likely predicted future. The results in Table III reveal that incorporating all uncertainties yields a high success rate but a lower episode reward compared to $p_{th}=0.15$. This indicates that while accounting for all uncertainties ensures task completion, it often results in deviations from the human trajectory and thus decreased rewards. Utilizing the most likely future leads to suboptimal performance in terms of both success and reward, highlighting the limitations of optimistic planning strategies.
TABLE III: Influence of probability threshold on planning performance
Threshold Success rate Episode reward Log divergence
0 0.98 44.85 10.49
0.15 0.98 46.62 10.27
$p_{\max}$ 0.94 44.28 10.39
V Conclusions
We develop an online and closed-loop behavior prediction model for POMDP planning in autonomous driving. Specifically, we propose an online recurrent neural belief update model and an option-based MCTS planner guided by Deep Q-learning. The belief update model can incorporate the ego vehicle’s intention to capture the closed-loop interactions between the AV and other agents. We introduce an online learning framework, which combines belief update learning and deep Q-learning to guide tree search. We validate our framework in simulated environments based on real-world driving scenarios. The experimental results indicate that our proposed online belief update model can significantly improve the temporal consistency and accuracy of probability estimations. Furthermore, DQN guidance considerably boosts the efficacy of the MCTS planner, resulting in improved decision-making performance. Future research could extend this framework to broader POMDP settings, such as noisy observations, by learning to update the belief states of agents to account for both actual physical states and future intentions. Moreover, the option generation process can be improved by learning from expert data and by making the options more flexible in terms of their length and termination criteria.
Acknowledgement
This work was supported by Berkeley DeepDrive. We would like to extend our gratitude to Yixiao Wang for his insights and support, as well as Rouhollah Jafari and Stefano Bonasera from General Motors for their constructive feedback.